### v2

# CaveAgent: Transforming LLMs into Stateful Runtime Operators

Abstract
Abstract:
LLM-based agents are increasingly capable of complex task execution, yet current agentic systems remain constrained by text-centric paradigms that struggle with long-horizon tasks due to fragile multi-turn dependencies and context drift. We present CaveAgent, a framework that shifts tool use from “LLM-as-Text-Generator” to “LLM-as-Runtime-Operator.” CaveAgent introduces a dual-stream architecture that inverts the conventional paradigm: rather than treating the LLM’s text context as the primary workspace with tools as auxiliary, CaveAgent elevates the persistent Python runtime as the central locus of state, with a lightweight semantic stream serving as its orchestrator. Beyond leveraging code generation to resolve interdependent sub-tasks (e.g., loops, conditionals) in a single step, CaveAgent introduces Stateful Runtime Management: it injects, manipulates, and retrieves complex Python objects (e.g., DataFrames, database connections) that persist across turns, unlike existing code-based approaches that remain text-bound. CaveAgent further provides a runtime-integrated skill management system that extends the Agent Skills open standard, enabling ecosystem interoperability through executable skill injections. This persistence mechanism serves as a high-fidelity external memory that reduces context drift in multi-turn interactions and preserves processed data for downstream applications without information loss. Evaluations on Tau2-bench, BFCL, and case studies across representative SOTA LLMs demonstrate consistent improvements: a 10.5% success rate gain on retail tasks, 28.4% reduction in total token consumption for multi-turn scenarios, and 59% token reduction on data-intensive tasks, enabling CaveAgent to handle data scales that cause context overflow in both JSON-based and code-based agents. The accessible runtime state further provides programmatically verifiable feedback, enabling automated evaluation and reward signal generation without human annotation and establishing a structural foundation for future research in Reinforcement Learning with Verifiable Rewards (RLVR).
[…]
1F4C5 Date: Feb 18, 2026 (v.2)
[…]
1F4BB Code: https://github.com/acodercat/cave-agent
[…]
1F4E7 Main Contact: Zhenglin Wan (vanzl@u.nus.edu), Jun Song (junsong@hkbu.edu.hk)
Figure 1: Town Simulation: a toy example for Stateful Runtime-Mediated Multi-Agent Collaboration.
1 Introduction
Large Language Models (LLMs) have demonstrated strong knowledge acquisition and reasoning capabilities across diverse natural language processing tasks. Building on these capabilities, tool-integrated reasoning (TIR) enables LLM agents to interact with external tools and APIs in a multi-turn manner [^1] In this paper, we use tool use and function calling interchangeably. (lu2023chameleon, shen2023hugginggpt, patil2024gorilla, qu2025tool), expanding their information access and solution space. This has extended LLM agents to various domains, including scientific discovery (boiko2023emergent, bran2024chemcrow), mathematical problem-solving (gao2023pal, chen2022program), Web GUI navigation (zhou2023webarena, yao2022webshop), and robotics (driess2023palme, brohan2023rt2).
Despite this progress, the conventional protocol for tool use requires LLMs to conform to predefined JSON schemas and generate structured JSON objects containing precise tool names and arguments (qin2023toolllm, openai2023gpt4). For example, to retrieve stock data, the model must synthesize a JSON string like {"tool": "get_stock", "params": {"ticker": "AAPL", "date": "today"}}, requiring exact adherence to syntax and field constraints. This imposes three architectural limitations: 1) Rigid Control Flow: each turn executes a single tool call (or parallel batch) and serializes output back to context, introducing latency for tasks requiring sequential orchestration (wu2023autogen, shen2023hugginggpt); 2) Statelessness: each call is an isolated transaction with no persistent state across turns, so intermediate results must be serialized back into text, causing token overhead for complex data structures and cascading error propagation (qiao2023taskweaver, kim2023language); and 3) Limited Composability: JSON schemas express flat function signatures and cannot natively represent loops, conditionals, or variable dependencies, forcing multi-step logic into error-prone multi-turn dialogue (kim2023language, wang2024executable).
While recent works attempt to address these issues with code-based tool use (wang2024executable, yang2024if), they predominantly adopt a process-oriented paradigm where the runtime state remains internalized and text-bound. This creates a “textualization bottleneck”: variables are accessible to external systems only through text output, requiring serialization into text strings (e.g., printing a DataFrame) to communicate with the user (wang2024executable, yao2022react). This prevents the direct input and output of structured, manipulatable objects, making it inefficient or impossible to handle complex non-textual data (e.g., large datasets, videos) (qiao2023taskweaver) and interact with downstream tasks. To address these limitations:
We aim to build a system that utilizes Python’s "everything is an object” philosophy to enable full Object-Oriented function calling and interaction, delegating context engineering to a persistent runtime and allowing the direct injection and retrieval of high-fidelity objects without serialization loss, thereby fully leveraging the strong code-generation capability of LLMs.
We present CaveAgent, an open-source framework that introduces the concept of Stateful Runtime Management for LLM agents, shifting code-based tool use from “process-oriented function calling” to persistent “object-oriented state manipulation.” CaveAgent operates on a dual-stream architecture with two distinct streams: a semantic stream for reasoning and a runtime stream for state management and code execution. This represents a fundamental inversion of the conventional paradigm: whereas existing agents treat the LLM’s semantic context as the primary workspace with external tools as auxiliary, CaveAgent elevates the persistent runtime as the primary locus of computation and state, with the semantic stream serving as a lightweight orchestrator that generates code to manipulate it.
By injecting complex data structures (e.g., graphs, DataFrames) directly into the runtime as persistent objects, CaveAgent achieves a form of context engineering: the agent manipulates high-fidelity data via concise variable references, decoupling storage from the limited context window. Any intermediate result (e.g., DataFrames, planning trees, or key metadata) can be stored in persistent variables that the agent actively retrieves for later use or downstream applications (code as action, state as memory). This reduces progressive context degradation in multi-turn interactions (catastrophic1, catastrophic2), enables context compression, and provides error-free recall through the runtime serving as an external memory.
The persistent environment also enables few-step resolution of complex logical dependencies by using code to interact with multiple interdependent tools, allowing the agent to compose workflows (e.g., data filtering followed by analysis) in a few turns rather than through error-prone multi-round function calling (multi-turn1, qin2023toolllm). The runtime’s transparency makes agent behavior fully verifiable, supporting checks on both intermediate programmatic states and final output objects of any data type, thereby enabling fine-grained reward signals for Reinforcement Learning.
Finally, CaveAgent supports artifact handoff without information loss by returning native Python objects rather than text representations, enabling direct use in downstream tasks such as UI rendering, visualization, and structured validation. The runtime can be serialized and reloaded, preserving the agent’s complete state across sessions. This transforms the LLM from an isolated text generator into a stateful, interoperable computational component within broader software ecosystems.
The function-calling paradigm in CaveAgent also extends beyond single-agent capabilities to enable Runtime-Mediated Multi-Agent Coordination, as shown in Figure 1 and Figure 2. Unlike conventional frameworks where agents coordinate via lossy text message passing (li2023camel, generative_agents), CaveAgent enables agents to interact through direct state manipulation. A supervisor agent can programmatically inject variables into a sub-agent’s runtime to alter its environment or task context without ambiguous natural language instructions. Multiple agents can also operate on a unified shared runtime, achieving implicit synchronization: when one agent modifies a shared object (e.g., updating a global “weather” entity in a town simulation), the change is immediately visible to all peers through direct reference. This transforms multi-agent collaboration from serialized dialogue into precise, verifiable state flow (details in Appendix E). We summarize our contributions as follows:
Figure 2: Key Advantages of CaveAgent
  • We introduce CaveAgent, a tool-use framework built on Stateful Runtime Management. CaveAgent shifts the paradigm from process-oriented function calling to persistent, object-oriented state management. It achieves context compression and context-grounded memory recall by delegating context engineering to a persistent runtime, eliminating the token overhead and precision loss of textual serialization while enabling few-step resolution of logically interdependent tasks. CaveAgent also provides runtime-integrated skill management that extends the Agent Skills open standard with executable injections for ecosystem interoperability.
  • The framework’s programmatic inspectability is a structural property of the architecture: runtime state is deterministically accessible at any point, providing automated evaluation and fine-grained reward signals for Reinforcement Learning with Verifiable Rewards (RLVR) without requiring subjective human annotation.
  • We evaluate CaveAgent on standard benchmarks (e.g., Tau2-Bench) and provide case studies across various domains including geospatial analysis (Appendix G.3) and AutoML multi-agent coordination (Appendix E.4). We also demonstrate the potential to extend the paradigm to Stateful Runtime-Mediated Multi-Agent Coordination with qualitative results, opening directions for future research.
2 CaveAgent: Stateful Runtime Management
2.1 Core Methodologies
In this section, we present the design of CaveAgent. As illustrated in Figure 3, CaveAgent adopts a dual-stream architecture, maintaining two synchronized streams throughout the interaction lifecycle: a Semantic Stream for lightweight reasoning, and a Runtime Stream for stateful execution, observation, and context engineering. Unlike conventional agents where the LLM’s text context serves as the primary workspace and external tools are called as auxiliary services, CaveAgent inverts this relationship: the persistent runtime becomes the central locus of data storage, computation, and state management, while the semantic stream is reduced to a lightweight controller that generates code to operate on the runtime.
We model the agent’s task as a sequential decision process over a horizon $T$. At each turn $t\in[1,T]$, the agent receives a query or observation $x_{t}$ and must produce a response $y_{t}$. Unlike traditional formulations where the entire state is re-serialized into $x_{t}$, we introduce a latent runtime state $\mathcal{S}_{t}$ (we call it "in-runtime context"). The system evolution is thus defined by:
$h_{t}$ $=\text{LLM}(x_{t},h_{t-1})\quad$ $(\text{Semantic Stream: Context History})$ (1)
$\mathcal{S}_{t}$ $=\text{Exec}(c_{t},\mathcal{S}_{t-1})\quad$ $(\text{Runtime Stream: Persistent Environment})$ (2)
where $h_{t}$ represents the semantic history (“in-prompt context”) and $c_{t}$ is the executable code generated by the agent. The key design choice is the decoupling of $h_{t}$ and $\mathcal{S}_{t}$: the semantic stream tracks intent and lightweight reasoning for code generation, while the runtime stream maintains all data and execution state via the code generated by the semantic stream.
The Runtime Stream:
The execution kernel of the runtime stream is a persistent Python kernel (an IPython interactive shell). We conceptualize each interaction turn $t$ not as an isolated API call, but as a cell execution in a virtual Jupyter notebook.
  • Persistent Namespace: The state $\mathcal{S}_{t}$ comprises the global namespace $\mathcal{N}_{t}$, containing all variables, functions, and imported modules. When the agent executes code $c_{t}$ (e.g., x = 5), the modification to $\mathcal{N}_{t}$ persists to $\mathcal{N}_{t+1}$. This allows subsequent turns to reference x directly without requiring the LLM to memorize or re-output its value.
  • Stateful Injection: Tools are not only described in text; they are injected into $\mathcal{N}_{0}$ as live Python objects. This allows the agent to interact with stateful objects via calling tools that modify the object’s internal state across turns.
Figure 3: Framework Overview
The runtime stream can also assign values to new variables during interaction and inject them into the persistent namespace (in-runtime context). This enables large context in complex tasks, such as large DataFrames, graphs, or other data structures, to be managed entirely by the Python runtime stream as stateful variables. Their values are preserved natively in persistent runtime memory without requiring repeated serialization into text, eliminating the risk of hallucination from lossy textual representations.
Code as Action, State as Memory  The agent stores key information (such as reasoning chains and intermediate data analysis results) as persistent variables in the runtime context, retaining only a lightweight description and reference in its in-prompt context. The runtime thus functions as an external memory, allowing the agent to retrieve stored data as native Python objects, achieving context compression and reducing progressive context degradation in multi-turn interactions. This property addresses persistent challenges in agentic tool use, specifically memory, dynamic decision-making, and long-horizon reasoning (bfcl).
A natural question arises: what distinguishes storing intermediate results in runtime variables from persisting them to files? We identify three key differences. First, type fidelity: runtime variables preserve native Python object types with full method interfaces (e.g., a DataFrame retains .groupby(), .merge() operations), whereas file-based storage requires serialization that may lose type information or fail entirely for non-serializable objects such as database connections, trained models with custom layers, or open file handles. Second, access latency: runtime variables enable zero-cost retrieval within the same memory space, while file I/O introduces disk latency and requires explicit read/write operations in generated code. Third, lifecycle management: runtime variables are automatically scoped to the agent session and garbage-collected appropriately, whereas file-based approaches require explicit cleanup logic to avoid accumulating temporary artifacts. That said, file-based persistence offers advantages for large-scale data exceeding memory capacity and for checkpointing across agent restarts. We note that CaveAgent also supports storing intermediate results in files, but storing them in runtime variables better respects the persistent runtime paradigm, yielding a more unified system.
Programmatic state retrieval enables the extraction of manipulated Python objects for direct use in downstream applications. Unlike conventional agents that produce text outputs requiring parsing and reconstruction, CaveAgent exposes native objects (DataFrames, class instances, arrays) with full type fidelity. This enables UI rendering via direct object binding, RL reward computation through programmatic state inspection, validation via unit test assertions against returned structures, and lossless object passing in multi-agent systems (Appendix E). The agent thus transforms the LLM from an isolated text generator into the operator of a stateful, interoperable computational component.
The Semantic Stream:
Parallel to the runtime stream, the semantic stream uses the LLM to generate code that manipulates the runtime. It is also responsible for:
  • Prompt Construction: Dynamically generating system instructions that describe the signatures of available tools in $\mathcal{N}_{t}$, without dumping their full state (which may be large) into the in-prompt context window.
  • Observation Shaping: Captures execution outputs and enforces a length constraint $\tau(\cdot)$ to prevent context explosion. This feedback mechanism guides the agent to interact with the persistent state efficiently, prioritizing concise and relevant information over verbose raw dumps in the in-prompt context $h_{t+1}$.
This dual-stream design addresses the “Context Explosion” problem: large data remains in the Runtime Stream ($\mathcal{S}_{t}$), while only high-level reasoning flows through the Semantic Stream ($h_{t}$). By making the runtime (not the LLM’s context window) the authoritative store of all data and intermediate state, CaveAgent avoids the token overhead and precision loss inherent in text-centric architectures. Unlike JSON-based function calling, which fails on inter-dependent tool calls (toolsandbox), CaveAgent dispatches interdependent tool chains in a few turns via executable code. Unlike code-based approaches with internalized runtimes, CaveAgent opens the runtime as a bidirectional interface for injecting and retrieving structured objects of any type. Algorithm 1 in Appendix A shows the iteration loop. We next describe the core designs of CaveAgent.
2.1.1 Variable and Function Injection
CaveAgent treats Python objects and functions as first-class citizens within the runtime. Each injectable entity is wrapped in a container that automatically extracts metadata: signatures, type hints, and docstrings for functions; names, types, and descriptions for variables. This metadata is aggregated into the system prompt as a lightweight “API reference,” while the actual objects are mapped directly into the execution engine’s namespace as global symbols. This design enables Object-Oriented Interaction: instead of stateless JSON calls (e.g., tool: "sort", args: {...}), the model invokes methods on stateful objects directly (e.g., processor.process(data)), chaining method calls and manipulating attributes naturally. The agent interacts via executable Python programs with native control flow (loops, conditionals) and stateful data passing, delivering final output as a native Python object rather than a textual approximation.
2.1.2 Dynamic Context Synchronization
The Semantic Stream is “blind” to the Runtime Stream by default. To inspect runtime state, the agent must explicitly generate code (e.g., print(df.head())), enforcing an Active Attention mechanism that selectively pulls only relevant slices into the token context. To prevent context explosion from verbose outputs, an Observation Shaping layer enforces a length constraint: when output exceeds $L_{\max}$, the system returns a structured error prompting the agent to use summary methods. This feedback loop guides efficient interaction with persistent state.
2.1.3 Security Check via Static Analysis
CaveAgent mitigates code execution risks via AST-based static analysis with modular policy rules: ImportRule (blocking unauthorized modules), FunctionRule (prohibiting dangerous calls like eval()), and AttributeRule (preventing sandbox bypass). Violations return structured errors to the semantic stream, enabling self-correction without breaking interaction continuity.
2.1.4 Runtime-Integrated Skill Management
Figure 4: CaveAgent’s Skill Management extends the Agent Skills open standard with runtime injection. Standard skills provide text prompts; CaveAgent skills additionally export functions, variables, and types into the persistent runtime.
CaveAgent extends the Agent Skills open standard [^2] https://agentskills.io, a portable specification adopted across the AI ecosystem including Claude Code, Gemini CLI, and Cursor. by introducing an injection.py module alongside the standard SKILL.md file. While standard skills provide text-based prompts that guide LLM behavior, CaveAgent skills additionally export Functions, Variables, and Type definitions that are injected directly into the persistent runtime upon activation. The framework employs progressive disclosure: skill metadata (name and description) is loaded at startup for routing decisions, while full instructions and runtime injections are loaded on-demand when the agent invokes activate_skill(). This bridges declarative skill definitions with CaveAgent’s stateful runtime paradigm: skills become executable capabilities, not merely textual instructions. As shown in Figure 4, domain expertise is packaged as both human-readable instructions for the language model and machine-executable artifacts for the runtime.
Table 1: Comparison of Standard Agent Skills vs. CaveAgent Skills.
Property Standard Skills CaveAgent Skills
Skill format SKILL.md SKILL.md + injection.py
LLM interaction Reads text prompts Operates on injected objects
Capability delivery Textual instructions Executable runtime artifacts
Data flow Text-in, text-out Object-in, object-out
3 Experiments
In this section, we validate CaveAgent by answering four questions:
  • [Q1.] Can CaveAgent perform on par with or surpass standard function-calling paradigms on widely-used benchmarks involving basic function-calling tasks? This is to showcase the basic function calling capabilities of CaveAgent.
  • [Q2.] Can CaveAgent successfully perform state management across multi-turns correctly and efficiently?
  • [Q3.] How token-efficient is CaveAgent compared to traditional JSON-based and Codeact style function calling?
  • [Q4.] How does CaveAgent adapt to complex scenarios that require manipulating complex data objects? This is to showcase CaveAgent’s unique advantages.
3.1 [Q1] Standard Function Calling Benchmarks
We evaluate on Tau2-bench (yao2024tau) and BFCL (bfcl) using six SOTA LLMs: DeepSeek-V3.2 (685B MoE), Qwen3 Coder (30B MoE), Kimi K2 0905 (1000B MoE), Claude Sonnet 4.5, GPT-5.1, and Gemini 3 Pro. For each model, we compare its native function-calling mechanism against CaveAgent, where the LLM serves solely as a text generation engine bypassing internal function-calling modules. Temperature settings follow official recommendations.
3.1.1 Results on Tau2-bench
Tau2-bench evaluates multi-turn tool use in realistic conversational scenarios (Airline and Retail domains), requiring agents to maintain consistency across turns. We follow the original evaluation protocols with DeepSeek V3 as user simulator, testing each model three times per domain. Since CaveAgent executes Python code, we employ runtime instrumentation with wrapper functions to capture and compare function invocations against ground truth, ensuring fair cross-paradigm evaluation.
Table 2: Performance comparison on the Tau2 Benchmark across different models and domains (3 Runs). Bold indicates CaveAgent outperforms Function Calling. The number inside each cell represents the success rate of each task (%).
Model Domain Function Calling CaveAgent
Run 1 Run 2 Run 3 Avg.↑ Run 1 Run 2 Run 3 Avg.↑
Open Source
DeepSeek-V3.2 (685B) Airline 56.0 56.0 54.0 55.3 62.0 60.0 58.0 60.0 (+4.7)
Retail 79.8 77.2 74.6 77.2 85.1 82.5 78.1 81.9 (+4.7)
Qwen3-Coder (30B) Airline 36.0 40.0 38.0 38.0 36.0 42.0 44.0 40.7 (+2.7)
Retail 41.2 43.0 39.5 41.2 51.8 54.4 57.9 54.7 (+13.5)
Kimi-K2-0905 (1000B) Airline 52.0 56.0 54.0 54.0 58.0 54.0 54.0 55.3 (+1.3)
Retail 62.3 60.5 59.6 60.8 69.3 72.8 71.9 71.3 (+10.5)
Closed Source
Claude Sonnet 4.5 Airline 56.0 54.0 62.0 57.3 56.0 52.0 62.0 56.7 (-0.6)
Retail 68.4 67.5 81.6 72.5 73.7 75.4 80.7 76.6 (+4.1)
GPT-5.1 Airline 50.0 58.0 50.0 52.7 58.0 56.0 54.0 56.0 (+3.3)
Retail 64.0 66.7 66.7 65.8 65.8 69.3 73.6 69.6 (+3.8)
Gemini 3 Pro Airline 64.0 62.0 58.0 61.3 68.0 68.0 68.0 68.0 (+6.7)
Retail 72.8 72.8 66.7 70.8 77.2 76.3 75.4 76.3 (+5.5)
Performance Analysis.
The results on Tau2-bench are summarized in Table 2. Key findings include:
(1). CaveAgent outperforms JSON-based function calling in 11 out of 12 settings across models from 30B to over 1000B parameters, with consistent improvements for DeepSeek-V3.2 (+5.3%) and Gemini 3 Pro (+6.1%), showing that offloading state management to a deterministic code runtime improves performance.
(2). Gains are amplified in state-intensive Retail scenarios, where complex transaction modifications require maintaining state consistency across turns. CaveAgent achieves double-digit gains for Qwen3 and Kimi K2, validating that Stateful Runtime Management reduces serialization-induced errors (detailed trajectory analysis in Appendix G.1).
(3). The code-specialized Qwen3-Coder (30B) exhibits the largest improvement (+13.5% in Retail), rivaling larger models. CaveAgent leverages the inherent coding proficiency of LLMs, allowing code-centric models to focus on logic generation rather than verbose context tracking.
3.1.2 Results on BFCL
To complement Tau2-bench’s multi-turn evaluation, we assess atomic function-calling precision on the Berkeley Function Calling Leaderboard (BFCL) (bfcl), which comprises $\sim$2,000 question-function-answer pairs across four categories of increasing complexity: simple, multiple, parallel, and parallel-multiple function calling. We use Executable Evaluation (functional correctness) by executing generated code and comparing results against ground truth. The summary is shown in Table 3 (detailed per-run results in Appendix Table 7).
Table 3: BFCL Benchmark Summary (avg. of 3 runs). Bold indicates CaveAgent outperforms Function Calling.
Model FC Avg.(%) CaveAgent Avg.(%) $\Delta$
DeepSeek-V3.2 (685B) 86.9 94.0 +7.1
DeepSeek-V3.2 (w/o prompt) 53.1 94.0 +40.9
Qwen3-Coder (30B) 89.8 94.4 +4.6
Kimi-K2-0905 (1000B) 89.2 94.7 +5.5
Claude Sonnet 4.5 94.4 94.4 0.0
GPT-5.1 89.6 88.9 -0.7
Gemini 3 Pro 94.3 94.3 0.0
Performance Analysis.
DeepSeek-V3.2 without explicit parallel-execution prompting achieves only 53.1% under the JSON paradigm due to its strong inductive bias toward sequential execution (liu2025deepseek). CaveAgent achieves 94.0% without any prompt intervention, as Python code naturally supports parallel execution via independent statements while preserving inter-tool dependency reasoning. The 30B Qwen3-Coder with CaveAgent (94.4%) outperforms the much larger GPT-5.1 (89.6%) and matches Claude Sonnet 4.5, demonstrating that CaveAgent leverages the coding proficiency of smaller LLMs. For SOTA models already at near-ceiling performance (Claude Sonnet 4.5, Gemini 3 Pro), gains are negligible since remaining errors stem from ambiguous queries rather than model incapacity. The advantages of our paradigm are most apparent in tasks requiring manipulation of complex data objects over long-horizon interactions, which we assess next.
3.2 [Q2] Case Study: Stateful Management
We design a benchmark targeting dimensions of state manipulation that existing benchmarks do not address, measuring an agent’s ability to read, modify, and persist variables across turns. A key design principle is programmatic validation: we directly inspect runtime state after execution against ground-truth expectations, rather than parsing text outputs. For each dimension, we curate test cases with linearly dependent queries and initial variable states (see Appendix D). Results are shown in Table 4.
Table 4: Stateful Management Benchmark Results (success rate %) across three evaluation dimensions. (1) For Type Proficiency, we designed 36, 36, and 42 cases for Simple, Object, and Scientific types, respectively. (2) For Multi-variable Stateful Management, we established 15 evaluation points for each variable-count tier. (3) For Multi-turn Stateful Management, we developed two scenarios, each consisting of 40 turns distributed across two conversations.
Type Proficiency (%) Multi-Variable (%) Multi-Turn (%)
Model Simple (36) Object (36) Sci. (42) Avg 5V (15) 10V (15) 15V (15) 20V (15) 25V (15) Avg Home (40) Fin. (40) Avg
DeepSeek-V3.2 100 100 100 100 100 100 100 100 100 100 100 100 100
Qwen3 Coder 100 94.4 95.2 96.5 94.4 100 80.0 80.0 100 90.9 77.5 85.0 81.3
Kimi K2 0905 100 100 100 100 100 100 100 100 100 100 90.0 100 95.0
Gemini 3 Pro 100 100 100 100 100 100 100 100 100 100 97.5 100 98.7
Type Proficiency
evaluates manipulation of Python primitives, user-defined class instances, and scientific types (DataFrames, ndarrays). Results yield uniformly high scores (96.5%–100%), validating that code-based manipulation of complex types is tractable for current LLMs.
Multi-Variable
tests how accuracy scales with 5–25 concurrent variables across five tiers (15 evaluation points each). Top models maintain 100% accuracy throughout, demonstrating that concurrent state management scales effectively within CaveAgent’s architecture.
Multi-Turn
assesses state persistence across 40-turn interactions in two scenarios: Smart Home (device state consistency) and Financial Account (numerical precision over multi-step operations). While DeepSeek-V3.2 maintains perfect accuracy, other models exhibit degradation on long-horizon state tracking. The consistently high accuracy across top models validates our thesis: when LLMs interact through code with persistent runtime state, reliable and verifiable agent behavior becomes achievable.
(a) Compare Execution steps and task success rates.
(b) Compare Prompt tokens and completion tokens.
Figure 5: Performance comparison between CaveAgent and traditional JSON-based Function Calling across three scenarios, focusing on execution steps, success rates, prompt token and completion token consumptions. Here, the steps means the number of turns needed for task completion, prompt tokens refers to the cumulative input tokens sent to the model across all turns (including system prompts, conversation history, and tool results), and completion tokens refers to the cumulative output tokens generated by the model (including reasoning, function calls, or code generation). The Func Call in right figure represents the abbrevation of Function Calling.
3.3 [Q3] Token Efficiency Study
We evaluate token efficiency across three domains (IoT, finance, e-commerce) with logically interdependent tool operations using DeepSeek V3.2. Figure 5 and Table 5 show results. CaveAgent achieves 28.4% lower total token consumption (504K vs. 704K) while improving success rate from 94.6% to 100%. The gain stems from resolving multiple dependencies in single code executions, reducing steps from 236 to 145 and prompt tokens by 32.7%. While CaveAgent consumes 36.3% more completion tokens (code is more verbose than JSON), prompt tokens dominate overall consumption and accumulate across turns.
Table 5: Summary of Improvements
Metric CaveAgent Function Calling Improvement
Total Steps 145 236 -38.6%
Prompt Tokens 444,679 660,588 -32.7%
Completion Tokens 59,440 43,600 +36.3%
Total Tokens 504,119 704,188 -28.4%
Avg. Success Rate 100% 94.6% +5.4 pp
3.4 [Q4] Case Study: Data-intensive Scenario
Table 6: Performance Comparison of three Function Calling paradigms across three task categories. CaveAgent (highlighted in green) achieves superior performance. Improvements relative to the best baseline are marked in parentheses.
Task Category Method Success Rate $\uparrow$ Prompt Tokens $\downarrow$ Compl. Tokens $\downarrow$ Total Tokens $\downarrow$
Data Query CaveAgent 100.0% (+20%) 118,901 4,584 123,485 (-51%)
CodeAct Style 80.0% 232,990 17,219 250,209
JSON-based FC 80.0% 278,239 16,413 294,652
Data Analysis CaveAgent 100.0% (Tie) 110,550 5,832 116,382 (-2%)
CodeAct Style 100.0% 112,990 6,232 119,222
JSON-based FC 10.0% 1,328,779 8,024 1,336,803
Visualization CaveAgent 90.0% (+50%) 374,855 30,250 405,105 (-39%)
CodeAct Style 40.0% 957,447 43,144 1,000,591
JSON-based FC 30.0% 644,778 17,899 662,677
We evaluate three architectures on a data-intensive benchmark comprising 30 tasks across data query, analysis, and visualization using stock market data (Apple and Google, 2020–2025). CodeAct Style disables CaveAgent’s variable injection/retrieval; JSON-based Function Calling operates without code execution. Results are shown in Table 6.
Results.
On Data Query, CaveAgent achieved 100% accuracy (123K tokens) by storing results in runtime variables, while both baselines failed at 80% due to context overflow from serializing large datasets. On Data Analysis, CaveAgent and CodeAct both achieved 100% with comparable tokens ($\sim$116–119K), but Function Calling managed only 10% (1.3M tokens) without code execution. On Visualization, CaveAgent achieved 90% (405K tokens) by retrieving chart data from runtime variables; CodeAct reached 40% (1M tokens) and Function Calling 30% (662K tokens). These results demonstrate that decoupling intermediate state from prompt context avoids the token accumulation causing context overflow in conventional architectures, with advantages growing with task complexity and data volume.
4 Conclusion
We present CaveAgent, a framework that transforms LLM tool use from stateless JSON function calling to persistent, object-oriented stateful runtime management. CaveAgent enables agents to maintain high-fidelity memory of complex objects and execute sophisticated logic via Python code. Experiments on Tau2-bench show that this approach outperforms SOTA baselines in multi-turn success rates (+10.5%) and token efficiency. Beyond performance gains, a key contribution is the programmatic verifiability enabled by CaveAgent’s architecture: because runtime state is deterministically inspectable, agent behavior can be evaluated automatically without human annotation, establishing a structural foundation for Reinforcement Learning with Verifiable Rewards and runtime-mediated multi-agent coordination. Qualitative case studies are provided in Appendix G.
Limitations.
CaveAgent requires tools to be Python-callable or wrapped as Python functions; tools with non-Python interfaces (e.g., native binaries, proprietary REST APIs) require manual wrapper creation. The persistent runtime consumes memory proportional to the complexity of stored objects; for extremely long sessions with large data artifacts, memory management becomes a concern. While agent behavior is in principle programmatically verifiable, designing comprehensive fine-grained benchmarks for stateful evaluation remains future work. Finally, the multi-agent coordination capabilities are demonstrated qualitatively; rigorous quantitative evaluation of runtime-mediated multi-agent systems is left for future investigation.
5 Background and Related Work
5.1 Tool Learning & Function Calling
The foundational approach to LLM tool use relies on a JSON-centric paradigm. The ReAct framework (yao2022react) generates actions as free-text strings requiring heuristic parsing, making it prone to hallucination and format errors (patil2024gorilla, liu2023agentbench). To address this, JSON-Schema function calling (patil2024gorilla, qin2023toolllm) constrains action generation to structured schemas, formalized by GPT-4 Function Calling and adopted by frameworks such as AutoGen (wu2023autogen). Constrained decoding methods like xGrammar (dong2024xgrammar) can further enforce syntactically valid JSON at low overhead. However, format correctness does not resolve deeper architectural limitations: JSON is a static interchange format lacking native control flow, its verbose syntax incurs token overhead (wang2024executable), and rigid schemas increase the probability of hallucination on complex nested structures (patil2024gorilla). The entire pipeline operates as a text-based serialization loop: the schema is flattened into the prompt, the LLM predicts a JSON payload, and middleware serializes results back into text, inheriting context explosion and error propagation (packer2023memgpt, wang2023mint).
5.2 Code as Action & Programmatic Reasoning
Recognizing these limitations, the “Code as Action” paradigm uses executable Python as a unified medium for reasoning and tool invocation. wang2024executable proposed CodeAct, reducing multi-turn overhead by up to 30% and improving task success rates by 20%. This shift leverages the Turing-complete nature of code to express loops, conditionals, and variable dependencies. The paradigm extends to domain-specific reasoning: ViperGPT (suris2023vipergpt) composes vision modules into executable subroutines, while Program of Thoughts (chen2022program) and PAL (gao2023pal) delegate arithmetic and symbolic logic to a Python interpreter. However, current code agents retain architectural limitations: CodeAct provides no explicit APIs for external object injection or retrieval, relying on a textualization bottleneck where intermediate states must be serialized via print and external data must be loaded through file I/O (wang2024executable). This makes it difficult to inject pre-existing Python objects (e.g., in-memory DataFrames, trained models) and risks context explosion with complex artifacts (liu2024lost, packer2023memgpt).
5.3 Context Management & Stateful Architectures
packer2023memgpt introduced MemGPT, an OS-inspired virtual context management system with tiered memory for long-horizon tasks. qiao2023taskweaver proposed TaskWeaver, a code-first framework preserving data structures across turns. However, existing approaches rely on RAG or textual summarization, inherently lossy methods that strip complex runtime objects of structural integrity and executable properties. CaveAgent uses Variable Injection to treat the Python runtime itself as high-fidelity external memory, allowing variables to persist in their native object form without re-tokenization overhead.
5.4 Multi-Agent Coordination
li2023camel proposed CAMEL for role-playing cooperation; qian2023chatdev introduced ChatDev with chat-chain workflows; hong2023metagpt developed MetaGPT encoding SOPs into prompts. These frameworks rely on text-based message passing, introducing serialization bottlenecks when transferring complex state between agents. CaveAgent enables Runtime-Mediated State Flow: agents collaborate by directly injecting and retrieving variables in a shared runtime, shifting coordination from “communication by talking” to “communication by shared state.”
Figure 6 illustrates the evolution of paradigm shifts in agentic tool use that motivates our work.
6 Acknowledgment
  • We thank Rui Zhou, a professional UI designer at Metasequoia Tech, for his assistance with the figure design in this paper.
  • We thank Qiuyang Mang, a Ph.D student in Computer Science at UC Berkeley, for the discussion about the core design of our framework.
Appendix
Appendix A Pseudo Code
Algorithm 1 shows the general workflow of CaveAgent.
Algorithm 1 CaveAgent Interaction Loop
1:Query $q$, Tools $\mathcal{T}$, Max Turns $T_{\max}$
[…]
4:for $t=1$ to $T_{\max}$ do
5:  Phase 1 (Reasoning): $R_{t}\leftarrow\text{LLM}(H_{t-1})$ $\triangleright$ Sample thought & code
6:  if $R_{t}$ contains code block $c_{t}$ then
7:   Phase 2 (Security): $V\leftarrow\textsc{ASTCheck}(c_{t},\Pi)$ $\triangleright$ Pre-exec validation
8:   if $V\neq\emptyset$ then
9:     $o_{t}\leftarrow\text{FormatError}(V)$
10:   else
11:     Phase 3 (Execution): $o_{t},\mathcal{S}_{t}\leftarrow\textsc{Run}(\mathcal{S}_{t-1},c_{t})$ $\triangleright$ Stateful update
12:     Phase 4 (Shaping): $o_{t}\leftarrow\textsc{Shape}(o_{t},L_{\max})$ $\triangleright$ Truncate & format
13:   end if
14:   $H_{t}\leftarrow H_{t-1}\cup\{(R_{t},o_{t})\}$ $\triangleright$ Sync observation to history
15:  else
16:   return $R_{t}$ $\triangleright$ Output final answer
17:  end if
18:end for
19:return "Max steps reached"
Figure 6: Evolution of Agentic Tool Use
Appendix B What Happens in Semantic Stream
The following sections detail the prompt templates used to instruct the Semantic Stream in CaveAgent. The system prompt is dynamically constructed by combining the Agent Identity, Context Information (functions, variables, types), and Instructions.
B.1 System Prompt Construction
The full system prompt is composed using the following template structure. The placeholders (e.g., {functions}) are populated at runtime with the specific tools and variables available in the current environment.
System Prompt Template {agent_identity} Current time: {current_time} You have access to: <functions> {functions}
</functions>
<variables> {variables}
</variables>
<types> {types}
</types>
Instructions: {instructions} {additional_context}
Below are the default values for the key components referenced in the template above.
Component: Agent Identity You are a tool-augmented agent specializing in Python programming that enables function-calling through LLM code generation. You have to leverage your coding capabilities to interact with tools through a Python runtime environment, allowing direct access to execution results and runtime state. The user will give you a task and you should solve it by writing Python code in the Python environment provided.
Component: Core Instructions 1. Carefully read and analyze the user’s input. 2. If the task requires Python code: - Generate appropriate Python code to address the user’s request. - Your code will then be executed in a Python environment, and the execution result will be returned to you as input for the next step. - During each intermediate step, you can use ’print()’ to save whatever important information you will then need in the following steps. - These print outputs will then be given to you as input for the next step. - Review the result and generate additional code as needed until the task is completed. 3. CRITICAL EXECUTION CONTEXT: You are operating in a persistent Jupyter-like environment where: - Each code block you write is executed in a new cell within the SAME continuous session - ALL variables, functions, and imports persist across cells automatically - You can directly reference any variable created in previous cells without using locals(), globals(), or any special access methods. 4. If the task doesn’t require Python code, provide a direct answer based on your knowledge. 5. Always provide your final answer in plain text, not as a code block. 6. You must not perform any calculations or operations yourself, even for simple tasks like sorting or addition. 7. Write your code in a {python_block_identifier} code block. In each step, write all your code in only one block. 8. Never predict, simulate, or fabricate code execution results. 9. To solve the task, you must plan forward to proceed in a series of steps, in a cycle of Thought and Code sequences.
B.2 Context Injection Format
Examples of how context is formatted for the LLM.
Example: Function Injection <functions> - function: buy_stock(symbol: str, quantity: int) -> Transaction description: Executes a stock purchase for the current portfolio. doc: xxArgs: xxxxsymbol: The ticker symbol of the stock (e.g., ’AAPL’). xxxxquantity: The number of shares to purchase. xxReturns: xxxxA Transaction object recording the details of the purchase. </functions>
Example: Variable Injection <variables> - name: portfolio type: Portfolio description: The user’s current investment portfolio object. - name: market_data type: DataFrame description: A pandas DataFrame containing historical price data. </variables>
Example: Type Schema Injection <types> Portfolio: xxdoc: Manages a collection of stock holdings and cash balance. xxmethods: xxxx- get_total_value() -> float xxxx- get_holdings() -> Dict[str, int] xxxx- add_cash(amount: float) -> None Transaction: xxdoc: An immutable record of a stock transaction. xxfields: xxxx- id: str xxxx- symbol: str xxxx- quantity: int xxxx- price_at_execution: float xxxx- timestamp: datetime </types>
B.3 Runtime Feedback Prompts
The agent operates in a closed feedback loop. After each code execution step, the runtime environment captures the output (stdout or errors) and constructs a new user message to guide the agent’s next action.
B.3.1 Standard Execution Output
This prompt is used when code executes successfully. It provides the standard output and explicitly reminds the agent that the variable state has been preserved.
Execution Output Template <execution_output> {execution_output}
</execution_output>
IMPORTANT CONTEXT REMINDER: - Based on this output, should we continue with more operations? - If the output includes an error, please review the error carefully and modify your code to fix the error if needed. - If yes, provide the next code block. If no, provide the final answer (not as a code block). - You are in the SAME Jupyter-like session. All variables from your previous code blocks are still available and can be accessed directly by name. - You DO NOT need to use locals(), globals(), or any special methods to access them. - Think of this exactly like working in Jupyter: when you create a variable in cell 1, you can simply use it by name in cell 2, 3, 4, etc.
B.3.2 Error Handling & Constraints
The system includes specific templates for handling edge cases, such as context window limits and security violations.
Output Length Exceeded: Used when the code generates excessive output (e.g., printing a massive DataFrame), prompting the agent to summarize instead.
Output Truncation Template The code execution generated output_length characters of output, which exceeds the maximum limit of max_length characters. Please modify your code to: 1. Avoid printing large datasets or lengthy content 2. Use summary statistics instead of full data (e.g., print shape, head(), describe() for dataframes) 3. Print only essential information needed for the task """
Security Violation: Used when the static analysis security checker blocks unsafe code (e.g., os.system).
Security Error Template <security_error> {error}
</security_error>
Code blocked for security reasons. Please modify your code to avoid this violation.
Appendix C What Happened in Runtime Stream
While the Semantic Stream governs reasoning and planning, the Runtime Stream serves as the execution engine and persistent memory. This stream operates as a dedicated Python kernel where data manipulation, tool invocation, and state transitions occur. The two streams follow a strict chronological topology, synchronized through interleaved exchange of code instructions and execution feedback.
C.1 Environment Initialization via Injection
The runtime lifecycle begins with Context Injection. Before the reasoning cycle starts, the user (or the system orchestration layer) initializes the runtime environment by injecting native Python objects directly into the global namespace.
  • Function Injection: Tool definitions are loaded as executable Python callables. Unlike RESTful API wrappers, these are native functions that can be inspected and invoked directly.
  • Variable Injection: Domain-specific data, such as DataFrames, graph structures, or class instances, are instantiated within the runtime stream’s memory.
This initialization phase populates the <functions> and <variables> blocks described in Section B.
C.2 The Interleaved Execution Paradigm
Once initialized, the workflow proceeds as a synchronized dialogue between the Semantic Stream (Reasoning) and the Runtime Stream (Execution). We conceptualize this as a dual-column timeline where actions are interleaved strictly in chronological order:
  1. 1.
    Semantic Turn (Left Cell): The LLM analyzes the current task and available context. It generates a Thought followed by a discrete Code Block (the instruction). This represents the input to the runtime.
  2. 2.
    Runtime Turn (Right Cell): The system extracts the code block and executes it within the persistent Python kernel. This execution constitutes the state transition $S_{t}\rightarrow S_{t+1}$. Crucially, this is not a stateless function call; it is a stateful operation where:
    • New variables defined in this cell are persisted in memory.
    • Existing objects (e.g., a list or a database connection) are mutated in place.
    • Side effects (e.g., saving a file) are realized immediately.
  3. 3.
    Feedback Loop: Upon completion of the Runtime Turn, the standard output (stdout), standard error (stderr), or the return value of the last expression is captured. This raw execution result is wrapped in the <execution_output> tags and injected back into the Semantic Stream, triggering the next Semantic Turn.
This mechanism ensures that the agent’s reasoning is always grounded in the current, actual state of the runtime environment.
C.3 Illustrative Case Study
To intuitively demonstrate the temporal synchronization and state dependency between the two streams, we present a concrete walkthrough in Figure 7. This example illustrates a toy data analysis task where the agent must filter a dataset and perform calculations on the result.
The workflow proceeds in a “zig-zag” pattern, alternating between reasoning (Left) and execution (Right):
  1. 1.
    Initialization ($T_{0}$): The user injects a pandas DataFrame named df. Note that the Semantic Stream only receives a lightweight pointer (variable name and documentation) instead of the whole data, while the Runtime Stream holds the actual heavy data object in memory.
  2. 2.
    Step 1 ($T_{1}\rightarrow T_{2}$): The agent generates code to filter the data. Crucially, the Runtime Stream does not return the full filtered dataset as text. Instead, it creates a new variable high_vol in the local scope and returns only a status update. This exemplifies our Stateful Management: the “result” of the tool use is a state change in memory, not a text string.
  3. 3.
    Step 2 ($T_{3}\rightarrow T_{4}$): The agent references the previously created variable high_vol to compute a statistic. This demonstrates Context Compression: the agent manipulates the data via variable references without ever consuming context tokens to “read” the full dataset.
Timeline of Interleaved Execution (Semantic vs. Runtime)
Time Semantic Stream (Reasoning & Code) Runtime Stream (State & Execution)
$T_{0}$ (Waiting for Input) [Initialization]
$\bullet$ df = pd.read_csv("market.csv") loaded into memory.
$\bullet$ <variables> metadata sent to Left Stream.
$T_{1}$ User: “Find stocks with volume > 1M.”
Agent Thought: I need to filter the dataframe df. I will save the result to a new variable.
Code:
high_vol = $df[$’volume’] > 1e6]
print(f"Filtered. Rows: {len(high_vol)}")
(Waiting for Instruction)
$T_{2}$ (Waiting for Feedback) [Execution]
1. Evaluates filter logic on df.
2.
New State: Variable high_vol created in RAM.
3.
Output: ‘‘Filtered. Rows: 42’’ sent to Left Stream.
$T_{3}$ Input: ‘‘Filtered. Rows: 42’’
Agent Thought: Now I need to calculate the average price of these high-volume stocks.
Code:
avg_price = high_vol[’price’].mean()
print(round(avg_price, 2))
(Waiting for Instruction)
$T_{4}$ (Waiting for Feedback) [Execution]
1. Accesses persistent object high_vol.
2. Computes mean.
3.
Output: 154.20 sent to Left Stream.
Figure 7: A visualization of the CaveAgent workflow. The process alternates between the Semantic Stream (generating instructions) and the Runtime Stream (executing and maintaining state). Note how the variable high_vol is maintained in the Runtime Stream ($T_{2}$) and accessed in the subsequent step ($T_{4}$) without re-loading or serialization, illustrating the efficiency of Stateful Runtime Management.
Viewing the runtime stream as a Jupyter notebook with multiple cells, where each cell corresponds to the execution at each time step, helps illustrate how states remain persistent across steps.
Appendix D Test Cases in Stateful Management Benchmark
In this section, we provide the examples of our test cases in Stateful Management Benchmark.
D.1 Type Proficiency Cases
The Type Proficiency category evaluates the agent’s ability to perform precise, state-aware manipulation of Python runtime elements. This section tests the agent’s working memory across three structural tiers: Simple Types (primitive types such as lists, dictionaries, and strings), Object Types (custom classes), and Scientific Types (high-dimensional complex data). Proficiency in these domains is a prerequisite for complex reasoning tasks.
D.1.1 Simple Types
Figure 8 shows the examples of our test cases of Simple types.
Simple Types
Turn User Query (Input) Immediate Validation (State Assertion)
Case: string_split_join
T1 “Set text to ’a,b,c’, split by comma, and rejoin with ’ | ’ as separator…” validate_str_split
$\bullet$ Assert text == "a | b | c".
T2 “Sort the parts of text alphabetically while keeping the ’ | ’ separator format.” validate_str_sort
$\bullet$ Assert text remains "a | b | c".
$\bullet$ Checks persistence of structure.
T3 “Reverse the order of parts in text but keep the ’ | ’ separator…” validate_str_reverse
$\bullet$ Assert text == "c | b | a".
Case: dict_nested
T1 “Change the math score to 90 in data[’scores’][’math’].” validate_dict_nested_update
$\bullet$ Assert data[’scores’][’math’] == 90.
T2 “The student just took a science test. Add a science score of 88 to data[’scores’].” validate_dict_nested_add
$\bullet$ Assert key ’science’ exists with value 88.
T3 “There was a curve on all tests. Add 5 points to every score in the scores dictionary.” validate_dict_increment
$\bullet$ Assert math == 95 (90+5).
$\bullet$ Assert science == 93 (88+5).
$\bullet$ Assert english == 95 (Initial 90+5).
Figure 8: Illustration of test cases of Simple Types. The results show the agent’s capability to manipulate any object types.
D.1.2 Object Types
Figure 9 shows the examples of our test cases of Object types.
Object Types
Turn User Query (Input) Immediate Validation (State Assertion)
Case: stack_advanced
T1 “Push ’A’, ’B’, ’C’, ’D’ in order.” validate_stack_multi_push
$\bullet$ Assert stack.size() == 4.
T2 “User wants to go back to first page. Pop until only 1 item remains, store count in result_num.” validate_stack_pop_until
$\bullet$ Assert stack.size() == 1.
$\bullet$ Assert result_num == 3 (Popped D,C,B).
T3 “Verify we’re at the right page. Peek at stack’s top and store in result_str.” validate_stack_peek_after
$\bullet$ Assert result_str == ’A’.
$\bullet$ Assert stack.size() == 1.
Case: cart_quantity
T1 “Add 3 Apples at \$10.00 each to cart with quantity.” validate_cart_qty_add
$\bullet$ Assert len(cart.items) == 1.
$\bullet$ Assert items[0][’quantity’] == 3.
T2 “Also add 2 Oranges at \$5.00 each…” validate_cart_qty_add2
$\bullet$ Assert len(cart.items) == 2.
T3 “Calculate total (price * quantity)… store in result_num.” validate_cart_qty_total
$\bullet$ Assert result_num == 40.0.
$\bullet$ Logic: $(3\times 10)+(2\times 5)$.
Figure 9: Illustration of test cases of Object Types. The results show the agent’s capability to manipulate custom class instances (Stack, ShoppingCart, Person) and verifying their internal attributes and method side-effects.
D.1.3 Scientific Types
Figure 10 shows the examples of our test cases of Scientific types.
Scientific Types
Turn User Query (Input) Immediate Validation (State Assertion)
Case: dataframe_merge (Relational Logic)
T1 “Merge df and df2 on product column. Store in result_df.” validate_df_merge
$\bullet$ Assert len(result_df) == 3.
$\bullet$ Assert column "supplier" exists.
T2 “Update result_df to keep only rows where supplier is ’SupA’.” validate_df_merge_filter
$\bullet$ Assert len(result_df) == 2.
$\bullet$ Logic: Keeps ’Phone’ and ’Shirt’.
T3 “Calculate the sum of prices in result_df. Store in result_value.” validate_df_merge_sum
$\bullet$ Assert result_value == 550.0.
$\bullet$ Logic: $500.0+50.0$.
Case: dataframe_pivot (Structure Reshaping)
T1 “Create pivot table from df_sales: region=rows, quarter=cols, sales=values.” validate_df_pivot
$\bullet$ Assert result_df.shape == (3, 2).
$\bullet$ Checks dimensions (3 regions, 2 quarters).
T2 “Calculate total sum of all sales…” validate_df_pivot_sum
$\bullet$ Assert result_value == 890.
$\bullet$ Verifies data integrity post-pivot.
T3 “Find which region has highest total sales (sum of Q1+Q2). Store sum…” validate_df_pivot_max_region
$\bullet$ Assert result_value == 380.
$\bullet$ Logic: South ($200+180$).
Case: ndarray_reshape (Tensor Manipulation)
T1 “Reshape array to shape (2, 4). Store in result_array.” validate_array_reshape
$\bullet$ Assert result_array.shape == (2, 4).
$\bullet$ Checks memory layout transformation.
T2 “Sum result_array along axis 1 (row sums).” validate_array_sum_axis
$\bullet$ Assert result equals [70, 48].
$\bullet$ Validates axis-wise reduction.
T3 “Calculate the total sum of result_array…” validate_array_total
$\bullet$ Assert result_value == 118.
$\bullet$ Logic: $70+48$.
Figure 10: Illustration of Scientific Types test cases. This benchmark challenges the agent with high-dimensionality operations, including relational merges (dataframe_merge), structural reshaping (dataframe_pivot), and tensor axis manipulation (ndarray_reshape), going beyond simple arithmetic.
D.2 Multi-variable Cases
Since there are 5 tiers of variable numbers, we select the variable number = 20 to demonstrate our test case since different variable number shares similar patterns of test cases. Figure 11 shows one example of test case where the agent is required to process 20 variables in 3 turns.
Multi-Variable Management: Tracking 20 Concurrent States (Full Context)
Turn Complex User Query (Full Text) State Verification (Partial View)
Case: startup_journey (20 Variables)
T1 “I’m documenting our startup TechStart. We’re in the Software industry, led by CEO Alice Johnson, headquartered in San Francisco. We have 50 employees, founded in $20$, 1 office, 2 products. Revenue is \$5M (5000000) with 10% profit margin (0.1), not public yet so no stock price or market cap. We’re profitable and hiring but not international yet. Departments: [’Engineering’, ’Sales’, ’Marketing’]. Locations: [’SF’]. Financials: funding 10000000, round ’Series A’. Contacts: email ’info@techstart.com’, phone ’555-0100’.” validate_startup_init
$\bullet$ employees $\to$ 50
$\bullet$ revenue $\to$ 5,000,000.0
$\bullet$ profit_margin $\to$ 0.1
$\bullet$ public $\to$ False
$\bullet$ stock_price $\to$ 0.0 (Initial)
T2 “Big growth update! Set employees to 150, offices to 3, products to 5. Set revenue to \$15M (15000000), profit_margin to 0.15. Set international to true. Append ’HR’ and ’Finance’ to departments. Append ’NYC’ and ’London’ to locations. Add ’valuation’: 100000000 to financials while keeping existing entries. Add ’support’: ’555-0200’ to contacts while keeping existing entries.” validate_startup_growth
$\bullet$ employees $\to$ 150
$\bullet$ depts $\to$ […, ’HR’, ’Finance’]
$\bullet$ financials $\to$ +{’valuation’: 100M}
$\bullet$ Assert founded_year == 2020 (Unchanged)
T3 “We’re going public! Append ’ Inc.’ to company_name. Set industry to ’Enterprise Software’. Set employees to 500, offices to 10, products to 10. Set revenue to \$50M (50000000), profit_margin to 0.2, stock_price to 25.0, market_cap to \$500M (500000000). Set public to true. Append ’Legal’ and ’IR’ to departments. Append ’Tokyo’ and ’Berlin’ to locations. Add ’ipo’: true to financials while keeping existing entries. Add ’ir’: ’irtechstart.com’ to contacts while keeping existing entries.” validate_startup_ipo
$\bullet$ public $\to$ True
$\bullet$ stock_price $\to$ 25.0
$\bullet$ company_name $\to$ "TechStart Inc."
$\bullet$ market_cap $\to$ 500,000,000.0
Figure 11: High-Dimensional State Management (Full Transcript). We present the raw input queries for the startup_journey case. The high information density requires the agent to parse and update over 10 distinct variables (Integers, Floats, Strings, Lists, Dictionaries) in a single turn (e.g., T3) without hallucination or omitting details.
D.3 Multi-turn Cases
These test cases evaluate the agent’s capability to process sequential instructions and maintain state precision over long-horizon scenarios. Unlike single-turn tasks where information is self-contained, these scenarios require the agent to maintain persistent memory of the system’s status, as subsequent queries depend on the outcome of previous actions. We categorize these multi-turn benchmarks into two domains: Smart Home Control and Financial Account Management.
D.3.1 Smart Home
In the Smart Home scenario, the agent acts as a central automation controller responsible for managing a suite of simulated IoT devices, including smart lighting, thermostats, motorized blinds, security cameras, and media players.
This benchmark specifically targets two advanced capabilities in stateful management:
  • Users frequently issue relative commands rather than absolute ones (e.g., “turn up the music more” or “dim the lights a bit”). To execute these correctly, the agent must recall the exact discrete level set in previous turns (e.g., incrementing volume from ’medium’ to ’high’) rather than resetting to a default value.
  • The agent must dynamically adjust device states based on simulated environmental contexts (e.g., “sunset”, “motion detected”) and complex user-defined conditions (e.g., “if the temperature drops below 10C, set heating to 22C”).
As illustrated in Figure 12, the weekend_party case spans a simulated 24-hour cycle. The agent must maintain a coherent environment state, transitioning from a quiet morning to a loud party and finally to a secure night mode, without drifting from the user’s cumulative intent.
Multi-Turn Scenario: Smart Home "Weekend Party" (Selected Turns)
Time / Turn User Query (Intent & Context) State Evolution & Validation
Turn 3
1:$0$ PM
“Party prep! Guests arriving soon. Adjust thermostat for comfort, set music to medium, open blinds fully, make lights bright.” validate_party_turn_3
$\bullet$ Music: OFF $\to$ 40% (Medium)
$\bullet$ Blinds: Closed $\to$ 100% (Full)
$\bullet$ Light: Dim $\to$ 80% (Bright)
Turn 5
4:$0$ PM
“Party mode! Full swing now. Turn up the music and make lights very bright. Verify camera is recording.” validate_party_turn_5
$\bullet$ Music: 50% $\to$ 60% (Party)
$\bullet$ Light: 80% $\to$ 90% (Very Bright)
$\bullet$ Camera: Assert status == Recording
Turn 7
7:$0$ PM
“Evening party. Close blinds completely, set mood lighting… turn up music more.” validate_party_turn_7
$\bullet$ Blinds: Partial $\to$ 0% (Closed)
$\bullet$ Music: 60% $\to$ 70% (Up More)
$\bullet$ Light: 90% $\to$ 60% (Mood)
Turn 10
10:$0$ PM
“Guests leaving. Lower music more, lock door, turn off bedroom light.” validate_party_turn_10
$\bullet$ Music: 80% $\to$ < 60% (Lowered)
$\bullet$ Door: Unlocked $\to$ Locked
$\bullet$ Bed Light: ON $\to$ OFF
Turn 17
Sun 10 AM
“Lazy morning… Finally getting up. Turn on bedroom light, open blinds, raise thermostat.” validate_party_turn_17
$\bullet$ Long-horizon consistency check
$\bullet$ Thermostat: Eco (18) $\to$ Comfort (21)
$\bullet$ Blinds: Closed $\to$ 70% (Open)
Figure 12: State persistence in long-horizon interactions. We visualize 5 key moments from the 20-turn weekend_party scenario. The agent must maintain a coherent environment state (lighting, temperature, security, audio) over a simulated 24-hour period. Crucially, it handles relative instructions (e.g., "turn up music", "lower music more") by tracking the exact discrete levels (e.g., Medium=40, Party=60) defined in the environment schema.
D.3.2 Financial Account
The Financial Account benchmark evaluates the agent’s capability to maintain strict numerical integrity and execute state-dependent logic within a banking ledger system. Unlike the relative adjustments in Smart Home, this domain demands exact integer arithmetic, where the agent must process a continuous stream of transactions, including deposits, interest applications, and loan amortizations, without cumulative drift.
This scenario imposes two constraints designed to stress-test the agent’s reasoning stability:
  • Operations require strict integer truncation (e.g., calculating $20\%$ of $1105$ as $221$, not $221.0$). Since the output of each turn (e.g., current balance) serves as the immutable basis for subsequent calculations (e.g., compound interest), a single arithmetic error in early turns triggers a cascading failure, rendering the entire subsequent interaction trajectory incorrect.
  • The agent must evaluate complex logic gates based on dynamic runtime states rather than static instructions. As demonstrated in the carol_debt_paydown case (Figure 13), queries often involve comparative functions (e.g., “pay the smaller of 15% of balance or 15% of loan”) or threshold checks (e.g., upgrading to ‘premium‘ status only if net worth becomes positive). This requires the agent to retrieve, compare, and act upon multiple variable states simultaneously before executing a transaction.
Multi-Turn Scenario: Financial Account "Carol’s Debt Paydown" (Numerical Precision)
Turn Conditional Query (Logic & Math) State Calculation & Assertions
T1 “Initialize account… Name ’Carol’, Balance 500, Status ’standard’, Interest 8% (Loan rate), Loan 2000.” validate_carol_turn_1
$\bullet$ Balance: 500
$\bullet$ Loan: 2000
$\bullet$ Status: ’standard’
T2 “Monthly loan interest due. Apply interest rate (8%) to loan balance and add to debt.” validate_carol_turn_2
$\bullet$ Interest = $2000\times 0.08=160$
$\bullet$ New Loan = $2000+160={\color[rgb]{0,0.390625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0.1953125}\texttt{2160}}$
T4 “Pay the smaller of 15% of balance or 15% of loan_balance. Subtract from both.”
(Context: T3 Paycheck +800 $\to$ Balance 1300)
validate_carol_turn_4
$\bullet$ [Logic] IF $\min(1300\times.15,2160\times.15)$
$\bullet$ Calc: $\min(195,324)=195$
$\bullet$ New Loan = $2160-195={\color[rgb]{0,0.390625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0.1953125}\texttt{1965}}$
T8 “Pay the larger of 40% of balance or 500 toward loan.”
(Context: Balance grew to 1574 after T7)
validate_carol_turn_8
$\bullet$ [Logic] Compare: $1574\times 0.4$ (629) vs 500
$\bullet$ Action: Pay 629
$\bullet$ Verify exact integer subtraction.
T14 “Check upgrade: IF loan_balance < balance, upgrade status to ’premium’.”
(Context: Loan reduced to 1172, Balance 1646)
validate_carol_turn_14
$\bullet$ [Logic] Condition: $1172<1646$ (True)
$\bullet$ Status $\to$ ’premium’
$\bullet$ Triggers T15 bonus paycheck.
T16 IF balance > loan_balance, pay off entire loan. Otherwise pay 75%…” validate_carol_turn_16
$\bullet$ [Logic] Action: Payoff Condition Met.
$\bullet$ Loan $\to$ 0.0
$\bullet$ Balance reduced by remaining debt.
Figure 13: Numerical precision and state-dependent reasoning. In the carol_debt_paydown scenario, the agent must perform exact integer arithmetic while navigating complex logic gates (e.g., Turn 4’s "smaller of", Turn 14’s "net worth check"). A single miscalculation in early turns (e.g., T2 interest) would cascade, causing failures in subsequent logic checks (e.g., failing the T16 payoff condition), thus rigorously testing long-horizon numerical stability.
Appendix E Stateful Runtime-Mediated Multi-Agent Coordination
The function-calling paradigm in CaveAgent introduces three key contributions for multi-agent coordination; Figure 1 illustrates an example. In this paper, we focus on qualitative analysis and provide case studies to facilitate understanding, leaving rigorous quantitative evaluation for future work. We introduce the high-level ideas below.
E.1 Meta-Agent Runtime Control
Sub-agents are injected as first-class objects into a meta-agent’s runtime, enabling the meta-agent to programmatically access and manipulate child agent states through generated code. Rather than following predefined communication protocols, the meta-agent dynamically sets variables in sub-agent runtimes, triggers execution, and retrieves results, enabling adaptive pipeline construction, iterative refinement loops, and conditional branching based on intermediate states.
E.2 State-Mediated Communication
Inter-agent data transfer bypasses message passing entirely. Agents communicate through direct runtime variable injection: the meta-agent retrieves objects from one agent’s runtime and injects them into another’s as native Python artifacts (DataFrames, trained models, statistical analyses), preserving type fidelity and method interfaces without serialization loss.
E.3 Shared-Runtime Synchronization
For peer-to-peer coordination, multiple agents can operate on a unified runtime instance, achieving implicit synchronization without explicit messaging. When one agent modifies a shared object, all peers perceive the change immediately through direct reference. New entities injected into the shared runtime become instantly discoverable, enabling collaborative manipulation of a unified world model with low coordination overhead.
How the town simulation demonstrates this capability. When the meta-agent modifies the weather state, all resident agents observe the change through direct attribute access; when a new location and manager are injected, existing agents can immediately query and interact with them.
Together, these patterns transform multi-agent systems from lossy text-based message exchange into typed, verifiable state flow, enabling automated validation of inter-agent handoffs and integration with downstream pipelines.
E.4 AutoML Training Loop
We demonstrate CaveAgent’s hierarchical agent coordination through an AutoML training loop where an orchestrator agent programmatically manages sub-agent runtimes (Figure 14). The orchestrator injects raw data into a feature engineering agent’s runtime via inject(), triggers execution, then retrieves the transformed DataFrame via retrieve() and injects it into a trainer agent’s runtime, all as native Python objects without serialization. After training, the orchestrator extracts evaluation metrics directly from the trainer’s runtime, validates against target requirements via a check_requirements() function, and injects performance feedback back into both sub-agents for the next iteration. Crucially, sub-agents themselves are injected as variables into the orchestrator’s runtime, enabling the orchestrator to dynamically access and manipulate their internal states through generated code. This iterative refinement loop continues until programmatic convergence criteria are met, demonstrating CaveAgent’s unique capability for hierarchical multi-agent coordination with typed, bidirectional state flow and automated convergence verification.
Figure 14: Multi-agent AutoML pipeline: an orchestrator coordinates feature engineering and training sub-agents through runtime variable injection. Native Python objects (DataFrames, trained models) flow losslessly between agents across iterative refinement cycles.
Appendix F Detailed BFCL Benchmark Results
Table F shows the detailed per-run results of BFCL benchmark.
Table 7: Detailed Performance comparison on BFCL Benchmark (3 Runs). Data is presented in score/total format. The Avg. columns indicate the average overall percentage across 3 runs. Bold indicates CaveAgent outperforms Function Calling. Simp. = Simple Function, Mult. = Multiple Function, Para. = Parallel Function, P-M. = Parallel Multiple Function, Ov. = Overall.
Model Run Function Calling CaveAgent
Simp. Mult. Para. P-M. Ov. Avg.(%) Simp. Mult. Para. P-M. Ov. Avg.(%)
Open Source
DeepSeek-V3.2 (685B) R1 354/400 183/200 175/200 159/200 871/1000 382/400 192/200 185/200 178/200 937/1000
R2 353/400 185/200 167/200 159/200 864/1000 386/400 193/200 184/200 178/200 941/1000
R3 360/400 185/200 173/200 154/200 872/1000 86.9 384/400 192/200 186/200 180/200 942/1000 94.0 (+7.1)
DeepSeek-V3.2 (w/o prompt) R1 312/400 162/200 33/200 26/200 533/1000 382/400 192/200 185/200 178/200 937/1000
R2 316/400 162/200 29/200 23/200 530/1000 386/400 193/200 184/200 178/200 941/1000
R3 314/400 161/200 35/200 21/200 531/1000 53.1 384/400 192/200 186/200 180/200 942/1000 94.0 (+40.9)
Qwen3-Coder (30B) R1 381/400 185/200 166/200 167/200 899/1000 386/400 191/200 187/200 180/200 944/1000
R2 381/400 185/200 166/200 167/200 899/1000 387/400 189/200 189/200 181/200 946/1000
R3 381/400 185/200 164/200 167/200 897/1000 89.8 386/400 190/200 189/200 178/200 943/1000 94.4 (+4.6)
Kimi-K2-0905 (1000B) R1 372/400 183/200 170/200 168/200 893/1000 387/400 191/200 186/200 187/200 951/1000
R2 368/400 181/200 167/200 171/200 887/1000 381/400 189/200 188/200 186/200 944/1000
R3 373/400 185/200 173/200 165/200 896/1000 89.2 379/400 191/200 188/200 187/200 945/1000 94.7 (+5.5)
Closed Source
Claude Sonnet 4.5 R1 387/400 189/200 184/200 183/200 943/1000 382/400 189/200 185/200 187/200 943/1000
R2 388/400 190/200 183/200 182/200 943/1000 384/400 189/200 185/200 186/200 944/1000
R3 387/400 190/200 184/200 184/200 945/1000 94.4 385/400 189/200 184/200 186/200 944/1000 94.4 ($\scriptstyle\sim$0.0)
GPT-5.1 R1 366/400 183/200 174/200 173/200 896/1000 367/400 186/200 172/200 176/200 901/1000
R2 367/400 186/200 173/200 169/200 895/1000 354/400 184/200 174/200 174/200 886/1000
R3 367/400 185/200 174/200 172/200 898/1000 89.6 356/400 180/200 170/200 175/200 881/1000 88.9 (-0.7)
Gemini 3 Pro R1 380/400 190/200 187/200 185/200 942/1000 382/400 191/200 184/200 186/200 943/1000
R2 380/400 192/200 188/200 183/200 943/1000 378/400 194/200 187/200 185/200 944/1000
R3 384/400 190/200 188/200 182/200 944/1000 94.3 380/400 194/200 184/200 185/200 943/1000 94.3 ($\scriptstyle\sim$0.0)
Appendix G Features
G.1 Case Analysis in Tau2-bench
To validate the architectural advantages of CaveAgent, we analyzed trajectory differences on the Tau2-bench retail benchmark. CaveAgent achieved a 72.8% success rate (83/114) compared to 62.3% (71/114) for the baseline JSON agent (Kimi K2 backbone), yielding a 10.5% improvement. We conducted a root cause analysis on the 24 tasks where CaveAgent succeeded but the baseline failed.
G.1.1 Failure Taxonomy of the Baseline
Baseline failures were categorized into five distinct patterns (Figure 15). The dominant failure mode (37.5%) was Missing Critical Action, where the agent retrieved necessary information but failed to execute the final operation (e.g., return, cancel). This was often coupled with Incomplete State Exploration (16.7%), where the agent heuristically queried subsets of data (e.g., checking only one recent order) rather than performing the exhaustive search required by the query.
05510101515202025253030353540404545 Wrong Order Selection Incomplete State Exploration Missing Calculation Wrong Arguments Missing Critical Action 16.716.7%16.716.7%20.820.8%2525%37.537.5%Percentage of Failure Cases (N=24)
Figure 15: Distribution of failure modes in baseline JSON agent trajectories. Note: Categories are non-exclusive as complex tasks may exhibit multiple failures.
G.1.2 Architectural Advantages: Loops and Conditionals
The analysis reveals that CaveAgent’s improvements stem from its ability to generate programming constructs, specifically loops (used in 92% of winning cases) and conditionals (83%), which resolve the semantic gaps inherent in single-step function calling.
Exhaustive State Exploration via Loops.
Tasks requiring global search (e.g., "return the order sent to Texas") baffled the baseline agent, which typically checked only 1–2 arbitrary orders. In contrast, CaveAgent generated for-loops to iterate through all user orders. For instance, in Task 26, the agent iterated through user.orders, checked order.address.state for "TX", and correctly identified the target order without hallucination.
Listing 1: Snippet from Task 26 showing exhaustive search.
# CaveAgent: Systematic iteration ensures no order is missed
for order_id in user_details.orders:
order = get_order_details(order_id)
if "TX" in order.address.state:
return_delivered_order_items(order_id, ...)
Complex Conditional Logic.
The baseline struggled with tasks involving fallback logic (e.g., "modify item, but if price > \$3000, cancel order"). In Task 90, the JSON agent ignored the price constraint and attempted modification regardless. CaveAgent successfully modeled this decision tree using explicit if/else blocks, checking variable states (variant.price) before execution.
Precise Attribute Reasoning.
While JSON agents rely on the LLM’s internal attention to compare values (often leading to errors like cancelling the wrong order in Task 59), CaveAgent offloads reasoning to the Python interpreter. By storing intermediate results (e.g., timestamps) in variables and using comparison functions (e.g., min()), CaveAgent ensured precise argument selection for actions requiring temporal or numerical comparisons.
G.2 Smart Home
Figure 16 illustrates the mechanistic advantage of CaveAgent through a toy smart-home example. The architecture separates the Semantic Stream (logic generation) from the Runtime Stream (state storage). This design enables two key capabilities absent in standard JSON agents:
  • State Persistence: Variables (e.g., Thermostat, Door) are initialized once and retain their state across multiple turns, eliminating the need to hallucinate or re-query context.
  • Control Flow Execution: The agent generates executable Python code with conditionals
    (e.g., if not door_lock.is_locked:), allowing for precise, context-dependent state transitions rather than blind API execution.
Figure 16: Demonstration of CaveAgent in a smart-home example: the Semantic Stream interacts with the Runtime Stream by generating code that manipulates stateful objects (variables) in the persistent runtime.
G.3 Geospatial Analysis
To illustrate CaveAgent’s advantages in domain-specific applications, we consider an urban planning scenario in which a user draws two arbitrary regions on an interactive map and queries the differences in population density and urban development between them. The interactive map converts user-drawn regions into GeoJSON polygon objects, variable-length coordinate arrays with high-precision floating-point pairs, which are injected directly into CaveAgent’s persistent runtime as first-class Python variables. Upon receiving the natural language query, the LLM generates a single code block that references these injected geometries to perform zonal statistics against WorldPop raster data and extract land use features from OpenStreetMap via osmnx, chaining multiple domain-specific operations through native variable passing. Under the conventional JSON function calling paradigm, the same task would require at least five sequential LLM turns, querying population and land use statistics separately for each region before synthesizing text-serialized results, while also confronting the challenge of encoding complex polygon geometries as JSON string parameters, which risks truncation and introduces serialization overhead. CaveAgent resolves the entire query in a single turn with lossless data flow: geometries maintain full numerical precision, intermediate results (e.g., GeoDataFrames) persist as manipulable runtime objects, and the LLM synthesizes the final response from deterministic execution output (Figure 17). This case study demonstrates CaveAgent’s suitability for scientific and analytical domains where computation involves complex non-serializable data structures and precision-sensitive results.
Figure 17: Geospatial Analysis: a user draws two regions on an interactive map. CaveAgent injects the GeoJSON polygons as Python variables and resolves the entire spatial query in a single turn with lossless data flow.